OpenStudy (anonymous):

1/(x-2) - 3x/(x-1) = (2x+1)/(x^2-3x+2)

OpenStudy (anonymous):

\[\frac{ 1 }{ x-2 }-\frac{ 3x }{ x-1 }=\frac{ 2x+1 }{ x ^{2}-3x+2 }\]

OpenStudy (anonymous):

I started by factoring the denominator on the right side: \[\frac{ 1 }{ x-2 }-\frac{ 3x }{ x-1 }= \frac{ 2x+1 }{ (x-2)(x-1) }\]

OpenStudy (anonymous):

And then used this as an LCD: \[\frac{ 1(x-1) }{ (x-2)(x-1) }-\frac{ 3x(x-2)}{ (x-2)(x-1) } = \frac{ 2x+1 }{ (x-2)(x-1) }\]

OpenStudy (anonymous):

I'm not sure what to do after this. It seems I'm messing up signs each time I try this equation. \[\frac{ x-1 }{ (x-2)(x-1) }-\frac{ 3x ^{2}-6}{ (x-2)(x-1) } = \frac{ 2x+1 }{ (x-2)(x-1) }\]

OpenStudy (anonymous):

The answer is x= 2/3, but I'm not sure how to get there.

OpenStudy (anonymous):

your last line is wrong when your multiplying through the LCD: I find it simpler to look at it as: \[\frac{1(x-1)-3x(x-2)}{(x-2)(x-1)}=\frac{2x+1}{(x-2)(x-1)}\] multiply both sides by (x-2)(x-1) to get: \[1(x-1)-3x(x-2)=2x+1\] simplifiy: \[x-1-3x^2-6x=2x+1\] \[-3x^2+5x-2=0\] Then use quadratic formula to solve: \[x=\frac{-b +\sqrt{ b^2-4ac}} {2a} \] I get 2 answers: 2 or \[-\frac{2}{3}\]

OpenStudy (anonymous):

however, we know 2 is not a solution, because the equation at 2 is undefined.